jigglypuff314 (jigglypuff314):

Hello :) I need a little help with arccos((x-2)/2) = arctan(___) for 2 < x < 4

OpenStudy (ranga):

Let arccos((x-2)/2) = y cos(y) = (x-2)/2 Find tan(y)

jigglypuff314 (jigglypuff314):

I"m not sure how I could do that? my teacher was trying to get us to do a triangle idea like 

OpenStudy (anonymous):

\[ y^2+(x-2)^2 = 2^2 \]

OpenStudy (ranga):

Find the length of the side opposite to the angle theta. tan(theta) = opposite / adjacent

OpenStudy (anonymous):

\[ \tan(\theta) = \frac{y}{x-2} \]

OpenStudy (anonymous):



jigglypuff314 (jigglypuff314):

mmm I got y^2 = -x^2 + 4x ? :)

OpenStudy (anonymous):

oh wait hmmm

OpenStudy (anonymous):

Yeah, you would have to take the square root.

jigglypuff314 (jigglypuff314):

ok :) so I get (sqrt(4x - x^2)) / (x-2) as the arctan?

OpenStudy (anonymous):

Yes

jigglypuff314 (jigglypuff314):

Thank you! :D <3

OpenStudy (anonymous):

the positive root might also be a solution.

OpenStudy (anonymous):

I mean negative root^

jigglypuff314 (jigglypuff314):

because of the given domain?

OpenStudy (anonymous):

because \(y\) was squared. Suppose \(y\) was negative.

jigglypuff314 (jigglypuff314):

ohhhhh okay ^_^